# MagicDrive3D: Controllable 3D Generation for Any-View Rendering in Street Scenes

### v3

@@ 4.2 Relative Pose Control for Video Generation
Given scene descriptions and a sequence of camera poses $\{\mathbf{S}_{t},\mathbf{P}_{c,t}\}$, our video generator is responsible for multi-view video generation. Although many previous art for street view generation achieve expressive visual effects, such as [9, 32, 30, 29], their formulations leave out a crucial requirement for 3D modeling. Specifically, the camera pose $\mathbf{P}_{c,t}$ is typically relative to the LiDAR coordinate of each frame. Thus, there is no precise control signal related to the ego trajectory, which significantly determines the geometric relationship between views of different $t$s.
@@
In our video generation model, we amend such precise control ability by adding the transformation between each frame to the first frame, i.e., $\mathbf{T}^{0}_{t}$. To properly encode such information, we adopt Fourier embedding with Multi-Layer Perception (MLP), and concatenate the embedding with the original embedding of $\mathbf{P}_{c,t}$, similar to [9]. As a result, our video generator provides better 3D consistency across frames, most importantly, making the camera poses to each view available in the same coordinate, i.e., $[\mathbf{R}^{0}_{c,t},\mathbf{t}^{0}_{c,t}]=\mathbf{T}^{0}_{t}[\mathbf{R}_{c,t},\mathbf{t}_{c,t}]$.
@@
4.3 Enhanced GS for Generated Content
@@
As introduced in Section 3, 3DGS is a flexible explicit representation for 3D reconstruction. Besides, the fast training and rendering speed of 3DGS make it highly suitable for reducing generation costs in our scene creation pipeline. However, similar to other 3D reconstruction methods, 3DGS necessitates high cross-view 3D consistency at the pixel level, which unavoidably magnifies the minute errors in the generated data into conspicuous artifacts. Therefore, we propose improvements for 3DGS from the perspectives of prior, modeling, and loss, enabling 3DGS to tolerate minor artifacts in the generated camera view from our video generation model, thereby becoming a potent tool for enhancing geometric consistency in rendering.
Figure 4: Illustration of the local dynamic from two successive generated frames of Front-Left (FL) camera. Even though our video generation model retains fine 3D consistency, minor discrepancies are inevitable. Our DGS can effectively reconstruct the scene with awareness of such discrepancy.
@@
Prior: Consistent Depth Prior. As essential geometry information, depth is extensively utilized in street scene reconstruction, such as the depth value from LiDAR or other depth sensors used in [34, 35]. However, for synthesized camera views, LiDAR/depth information is unavailable (also making these methods not suitable for our scenario). Therefore, we propose to use pre-trained monocular depth estimator [3] to infer depth information.
@@
While monocular depth estimation is separate for each camera view, proper scale $s_{c,t}$ and offset $b_{c,t}$ parameters should be estimated to align them for a single scene [40], as in Figure 3(a). To this end, we first apply the Point Cloud (PCD) from Structure of Motion (SfM) [24, 23], shown in Figure 3(b). However, such PCD is too sparse to accurately restore $(s_{c,t},b_{c,t})$ for any views. To bridge the final gap, secondly, we propose further optimizing the $(s_{c,t},b_{c,t})$ using the GS loss, as in Figure 3(c). Specifically, we replace the optimization for Guassian centers $\boldsymbol{\mu}_{i}$ with $(s_{c,t},b_{c,t})$. After the optimization, we initialize $\boldsymbol{\mu}_{i}$ with points from depth values. Since GS algorithm is sensitive to accurate point initialization [12, 8], our method provides useful prior to reconstructing in this sparse view scenario.
@@
Modeling: Deformable Gaussians for Local Dynamic. Despite the 3D geometric consistency provided by our video generation model, there are inevitably pixel-level disparities in some object details, as shown in Figure 4. The strict consistency assumption of 3DGS may amplify these minor errors, resulting in floater artifacts. To mitigate the impact of these errors, we propose Deformable Gaussian Splitting (DGS), which, reduces the requirement for temporal consistency between frames, thereby ensuring the reconstruction quality of the generated views.
@@
Specifically, as shown in Figure 4, we pick the center frame $t=t_{C}$ as the canonical space and enforce all Gaussians in this space. Hence, we allocate a set of offsets to each Gaussian, $\boldsymbol{\mu}^{o}_{p}(t)\in\mathbb{R}^{3}$, where $t\in[1,\dots,T]$ and $\boldsymbol{\mu}^{o}_{p}(t_{C})\equiv\mathbf{0}$. Note that, different camera views from the same $t$ share the same $\boldsymbol{\mu}^{o}_{p}(t)$ for each Gaussian, and we apply regularization on them to keep the dynamic in local, as shown in Equation 2:
@@
Consequently, $\boldsymbol{\mu}^{o}_{p}(t)$ can manage the local dynamics driven by pixel-level disparities, while $\boldsymbol{\mu}$ focuses on the global geometric correlations. It ensures the quality of scene reconstruction by leveraging consistent parts across different viewpoints, simultaneously eliminating artifacts. Besides, with the analytical gradient w.r.t. $SE(3)$ pose of cameras [17], we also make the camera pose optimizable in the final few steps of GS iterations, which helps to mitigate the local dynamic from camera poses.
Algorithm 1 Enhanced Deformable GS
1:camera views $\{\mathcal{I}_{i}\}$, camera parameters $\{\mathbf{P}^{0}_{i}\}$, monocular depth $\{\mathcal{D}_{i}\}$, optimization steps for depth $s_{D}$, camera pose $s_{\text{C}}$, and GS $s_{\text{GS}}$
2:DGS of the scene $\{\boldsymbol{\mu}_{p},\boldsymbol{\mu}^{o}_{p},\boldsymbol{\Sigma}_{p},\text{SH}_{p}\}$, and optimized camera pose $\{\mathbf{P}^{0}_{i}\}$
3:$\mathcal{P}_{SfM}=$ PCD from SfM
4:Optimize $(s_{c,t},b_{c,t})$ with $\mathcal{P}_{\text{SfM}}$ for each $\{c,t\}$
5:Random initialize AEs $\{\mathbf{e}_{i}\}$
6:for step in $1,\dots,s_{D}$ do
7:    Random pick one view $\mathcal{I}_{i}$
8:    $\mathcal{L}=\mathcal{L}_{\text{AEGS}}(\mathcal{I}_{i},\mathcal{I}^{r}_{i},\mathbf{e}_{i})$
9:    Update $(s,b),\mathbf{e}_{i},\boldsymbol{\Sigma},\text{SH}$ with $\nabla\mathcal{L}$
10:end for
11:Initialize $\boldsymbol{\mu}$ with $(s,b)$ and $\mathcal{D}$
12:for step in $s_{D},\dots,s_{\text{GS}}$ do
13:    Random pick one view $\mathcal{I}_{i}$ and get its $t$
14:    $\mathcal{L}=\mathcal{L}_{\text{DGS}}(\mathcal{I}_{i},\mathcal{I}^{r}_{i},\mathbf{e}_{i},\boldsymbol{\mu}^{o}(t))$
15:    Update $\boldsymbol{\mu},\boldsymbol{\mu}^{o}(t),\mathbf{e}_{i},\boldsymbol{\Sigma},\text{SH}$ with $\nabla\mathcal{L}$
16:    if step $>s_{\text{C}}$ then
17:         Update $\mathbf{P}^{0}_{i}$ with $\nabla\mathcal{L}$
18:    end if
19:end for
Figure 5: Qualitative comparison with NF-LDM [13] and original 3DGS (on the same generated video as ours). Panoramas for GS are transformed and stitched from perspective cameras with $90^{\circ}$ FOV. Views in the last row are rendered with unseen camera rigs of nuScenes.
@@
Loss: Aligning Exposure with Appearance Modeling. Typical street view datasets are collected with multiple cameras, which capture views independently through auto-exposure and auto-white-balance [4]. Since the video generation is optimized to match the original data distribution, the differences from different cameras also exist in the generated data. The appearance differences are well-known issues for in-the-wild reconstruction [16]. In this paper, we propose a dedicated appearance modeling technique for GS representation.
@@
We hypothesize that the disparity between different views can be represented by affine transformations $\mathbf{A}_{i}(\cdot)$ for $i$-th camera view. An Appearance Embedding (AE) map $\mathbf{e}_{i}\in\mathbb{R}^{w_{e}\times h_{e}\times c_{e}}$ is allocated for each view, and a Convolutional Neural Network (CNN) is utilized to approximate this transformation matrix $w_{\mathbf{A}}\in\mathbb{R}^{w\times h\times 3}$ (Appendix F contains more details). The final computation of the pixel-wise $\ell_{1}$ loss is conducted using the transformed image. Therefore, our final loss for DGS is as Equation 3, where $\lambda_{{\text{reg}_{o}}}$ is the hyper-parameter for offset regularization.
$\begin{split}\mathcal{L}_{\text{DGS}}&=\mathcal{L}_{\text{AEGS}}+\lambda_{{\text{reg}_{o}}}\mathcal{L}_{{\text{reg}_{o}}}\\ &=(1-\lambda)\mathcal{L}_{1}(\mathbf{A}_{i}(\mathcal{I}^{r}_{i}),\mathcal{I}_{i})\\ &+\lambda\mathcal{L}_{\text{D-SSIM}}(\mathcal{I}^{r}_{i},\mathcal{I}_{i})+\lambda_{{\text{reg}_{o}}}\mathcal{L}_{{\text{reg}_{o}}}\end{split}$ (3)
Optimization Flow. We demonstrate the overall optimization flow of the proposed DGS in Algorithm 1. Line 2 is the first optimization of monocular depths. Lines 4-8 refer to the second optimization of the monocular depths. Lines 10-16 are the main loop for DGS reconstruction, where we consider temporal offsets on Gaussians, camera pose optimization for local dynamics, and AEs for appearance discrepancies among views.
@@ 5.1 Experimental Setup
Dataset. We test our MagicDrive3D using the nuScenes dataset [4], which is commonly used for generating and reconstructing street views [9, 32, 30, 34]. The official configuration is followed, using 700 street-view video clips of approximately 20s each for training and another 150 clips for validation. For semantics in control signals, we follow [9], using 10 object classes and 8 road classes.
  name #test #train camera poses
  360 6 90 [Uncaptioned image]
vary-t 12 84 [Uncaptioned image]
Table 1: Two settings for reconstruction quality evaluation. Testing views are in green while training views are in red.
Metrics and Settings. MagicDrive3D is primarily evaluated using the Fréchet Inception Distance (FID) by rendering novel views unseen in the dataset and comparing their FID with real images. In addition, the method’s video generation ability is evaluated using Fréchet Video Distance (FVD), and its reconstruction performance is assessed using L1, PSNR, SSIM [31], and LPIPS [37]. For reconstruction evaluation, two testing scenarios are employed: 360, where all six views from $t=9$ are reserved for testing the reconstruction in the canonical space; and vary-t, where one view is randomly sampled from different $t$ to assess long-range reconstruction ability through $t$ in the canonical space (as shown in Table 1). More implementation details can be found in Appendix A.
Methods FVD FID (seen) FID (novel)
MagicDrive[9] 177.26 20.92 N/A
Ours (video gen.) 164.72 20.67 N/A
3DGS N/A 45.07 145.72
Ours (scene gen.) N/A 23.99 34.45
Table 2: Generation quality evaluation. We adopt all validation scenes from nuScenes. All generated views are used for reconstruction. Novel views are camera poses different from nuScenes.
@@
train view vary-t 3DGS 0.0189 30.1191 0.9261 0.1259
@@
test view vary-t 3DGS 0.0890 17.9879 0.4378 0.4648
@@
Table 3: Reconstruction quality evaluation. We randomly sample 100 scenes from the nuScenes val. set for evaluation. “cc” refers to color correction from [1]. Although 3DGS does not consider appearance differences, we apply “cc” to it for fair comparisons.
5.2 Main Results
Generation Quality. As shown in Table 2, the evaluation of generation quality involves two aspects. Firstly, the quality of video generation is assessed using extended 16-frame MagicDrive [9] as a baseline. Despite minor improvement in single-frame quality based on FID, MagicDrive3D substantially enhances video quality (as evidenced by FVD), demonstrating the efficacy of the proposed relative camera pose embedding in enhancing temporal consistency. Secondly, the image quality of renderings from the generated scene is evaluated using FID. We also include qualitative comparisons in Figure 5. Compared to 3DGS, our enhanced DGS significantly enhances visual quality in reconstructing contents, particularly in unseen novel views. More qualitative comparison can be found in Appendix E.
Reconstruction Quality. Our enhanced DGS, as a reconstruction method, is further evaluated by comparing renderings with ground truth (GT) images. Here, the generated views from the video generator are treated as GT. We employ two settings per Table 1, with results displayed in Table 3. As per all metrics, our enhanced DGS not only improves reconstruction quality for training views but also drastically enhances quality for testing views, compared to 3DGS. We include comparison and discussion with 4DGS [33] in Appendix C.
@@ 5.2 Main Results
Table 4: Ablation study on enhanced DGS. We adopt the same settings as in Table 3, where 100 scenes from the nuScenes validation set are adopted.
(a) Original view.
(b) Move forward 0.1m.
(c) Move forward 1m.
(d) Move forward 1.6m.
Figure 6: Application on rendering object-level dynamic. After scene generation, we can segment and move the vehicle (the one on the right) in 3D to render a dynamic object.
Controllability. MagicDrive3D accepts 3D bounding boxes, BEV map, and text as control signals, each of which possesses the capacity to independently manipulate the scene. To show such controllability, we edit a scene from the nuScenes validation set, as presented in Figure 7. Clearly, MagicDrive3D can effectively alter the generation of the scene to align with various control signals while maintaining 3D geometric consistency.
Setting Method no rig depth+0.5m pitch-5 yaw+5 yaw-5
vehicle only real data 17.14 16.63 15.50 16.99 15.94
road only real data 54.94 54.56 53.82 54.20 53.67

### v4

@@ 4.2 Video Generation with Relative Pose Control
Given scene descriptions and a sequence of camera poses $\{\mathbf{S}_{t},\mathbf{P}_{c,t}\}$, our video generator is responsible for multi-view video generation. Although many previous arts for street view generation achieve expressive visual effects (e.g.,  [11, 36, 32, 31]), their formulations leave out a crucial requirement for 3D modeling. Specifically, the camera pose $\mathbf{P}_{c,t}$ they use is typically relative to the LiDAR coordinate of each frame. Thus, there is no precise control signal related to the ego trajectory, which significantly determines the geometric relationship between views of different $t$s.
@@
In our video generation model, we amend such precise control ability by adding the transformation between each frame to the first frame, i.e., $\mathbf{T}^{0}_{t}$. To properly encode such information, we adopt Fourier embedding with Multi-Layer Perception (MLP), and concatenate the embedding with the original embedding of $\mathbf{P}_{c,t}$, similar to Gao et al. [11]. As a result, our video generator provides better temporal quality across frames, most importantly, making the camera poses to each view available in the same coordinate, i.e., $[\mathbf{R}^{0}_{c,t},\mathbf{t}^{0}_{c,t}]=\mathbf{T}^{0}_{t}[\mathbf{R}_{c,t},\mathbf{t}_{c,t}]$.
@@
4.3 Enhanced FTGS for 3DGS Generation
@@
As introduced in Section 3, 3DGS is a flexible explicit representation for 3D scenes. Besides, the fast training and rendering speed of 3DGS makes it highly suitable for reducing generation costs in our scene creation pipeline. However, similar to other 3D reconstruction methods, 3DGS necessitates high cross-view 3D consistency at the pixel level, which unavoidably magnifies the minor artifacts in the generated data into conspicuous errors. Therefore, we propose improvements for 3DGS from the perspectives of prior, modeling, and loss, enabling 3DGS to tolerate minor artifacts in the generated camera view from our video generation model, thereby becoming a potent tool for enhancing geometric consistency in rendering.
@@
Prior: Consistent Depth Prior. As essential geometry information, depth is extensively utilized in street scene reconstruction, such as the depth value from LiDAR or other depth sensors used by Yan et al. [39]. However, for synthesized camera views, LiDAR/depth information is unavailable (also making these methods not suitable for our scenario). Therefore, we propose to use a pre-trained monocular depth estimator [3] to infer depth information.
@@
While monocular depth estimation is separate for each camera view, proper scale $s_{c,t}$ and offset $b_{c,t}$ parameters should be estimated to align them for a single scene [47], as in Figure 4(a). To this end, we first apply the Point Cloud (PCD) from Structure of Motion (SfM) [27, 26] for estimation, shown in Figure 4(b). However, such PCD is too sparse to accurately restore $(s_{c,t},b_{c,t})$ for any views. To bridge the gap, secondly, we propose further optimizing the $(s_{c,t},b_{c,t})$ using the GS loss, as in Figure 4(c). Specifically, we replace the optimization for Guassian centers $\boldsymbol{\mu}_{i}$ with $(s_{c,t},b_{c,t})$. After the optimization, we reset $\boldsymbol{\mu}_{i}$ with points from depth values. Since GS algorithm is sensitive to accurate point initialization [13, 9], our method provides a useful prior for 3DGS generation in this sparse view scenario.
Figure 5: Qualitative comparison with NF-LDM [14]. Our method can generate higher quality 3D scenes while maintaining better object geometry, with stronger controllability compared to NF-LDM (see the significant object deformation in NF-LDM). Panoramas for GS are transformed and stitched from perspective cameras with $90^{\circ}$ FOV. Views in the last row are rendered with unseen camera rigs of nuScenes.
@@
Modeling: Fault-Tolerant Gaussians for Local Inconsistency. Despite the 3D geometric consistency provided by our video generation model, there are inevitably pixel-level disparities in some object details, as shown in Figure 3. Even worse, the strict consistency assumption of 3DGS [13] (and all other reconstruction algorithms, e.g., [48, 38, 37]) can amplify these minor errors, resulting in floater artifacts. To mitigate the impact of these errors, we propose Fault-Tolerant Gaussian Splitting (FTGS), which reduces the requirement for temporal consistency between frames, thereby ensuring the 3DGS generation quality in the canonical space from the generated views.
@@
Specifically, as shown in Figure 3, we pick the center frame $t=t_{C}$ as the canonical space and enforce all Gaussians in this space. Then, we allocate a set of offsets to each Gaussian, $\boldsymbol{\mu}^{o}_{p}(t)\in\mathbb{R}^{3}$, where $t\in[1,\dots,T]$ and $\boldsymbol{\mu}^{o}_{p}(t_{C})\equiv\mathbf{0}$. Note that, different camera views from the same $t$ share the same $\boldsymbol{\mu}^{o}_{p}(t)$ for each Gaussian, and we apply regularization on them to keep the offsets in local, as shown in Equation 2:
@@
Consequently, $\boldsymbol{\mu}^{o}_{p}(t)$ can manage the local inconsistency driven by pixel-level disparities, while $\boldsymbol{\mu}$ focuses on the global geometric correlations. It ensures the quality of scene 3DGS generation by leveraging consistent parts across different viewpoints, simultaneously eliminating artifacts. Besides, with the analytical gradient w.r.t. $SE(3)$ pose of cameras [20], we make the camera pose optimizable in the final few steps of GS iterations, which helps to mitigate the local inconsistency from camera poses.
@@
Loss: Aligning Exposure with Appearance Modeling. Typical street view datasets are collected using multiple cameras, each operating independently with auto-exposure and auto-white-balance [4]. Appearance modeling addresses issues from such in-the-wild reconstructions [19]. Similarly, video generation is optimized to match the real data distribution, so differences between cameras also appear in generations. We propose a dedicated appearance modeling technique for GS representation as a solution.
@@
We hypothesize that the disparity between different views can be represented by affine transformations $\mathbf{A}_{i}(\cdot)$ for $i$-th camera view. An Appearance Embedding (AE) map $\mathbf{e}_{i}\in\mathbb{R}^{w_{e}\times h_{e}\times c_{e}}$ is allocated for each view, and a Convolutional Neural Network (CNN) is utilized to approximate this transformation matrix $w_{\mathbf{A}}\in\mathbb{R}^{w\times h\times 3}$ (Appendix H contains more details). The final computation of the pixel-wise $\ell_{1}$ loss is conducted using the transformed image.
Therefore, our final loss for FTGS is as Equation 3, where $\lambda_{{\text{reg}_{o}}}$ is the hyper-parameter for offset regularization.
$\mathcal{L}_{\text{FTGS}}$ $=\mathcal{L}_{\text{AEGS}}+\lambda_{{\text{reg}_{o}}}\mathcal{L}_{{\text{reg}_{o}}}$ (3)
$=(1-\lambda)\mathcal{L}_{1}(\mathbf{A}_{i}(\mathcal{I}^{r}_{i}),\mathcal{I}_{i})+\lambda\mathcal{L}_{\text{D-SSIM}}(\mathcal{I}^{r}_{i},\mathcal{I}_{i})+\lambda_{{\text{reg}_{o}}}\mathcal{L}_{{\text{reg}_{o}}}$
Appendix C holds a full description of the algorithm.
Methods FVD FID (seen) FID (novel)
MagicDrive 177.26 20.92 N/A
MagicDrive3D (video gen.) 164.72 20.67 N/A
3DGS N/A 45.07 145.72
MagicDrive3D (scene gen.) N/A 23.99 34.45
Table 2: Generation quality evaluation. We adopt all val. scenes from nuScenes. All generated views are used for 3DGS generation. Novel views using camera poses different from nuScenes.
@@ 5.1 Experimental Setup
Dataset. We test MagicDrive3D using nuScenes dataset [4], which is commonly used for generating and reconstructing street views [11, 36, 32, 38]. The official configuration is followed (700 street-view video clips of $\sim$20s each for training and another 150 clips for validation). For semantics in control signals, we follow Gao et al. [11], using 10 object classes and 8 road classes.
Metrics and Settings. There is no direct metric for 3D scene generation. Following Kim et al. [14], we primarily evaluate MagicDrive3D using the Fréchet Inception Distance (FID) by rendering novel views unseen in the dataset and comparing their FID with real images.
To further demonstrate the effectiveness of each component within our framework, we conducted ablated studies on the improvements in video generation capabilities and the enhancements in content reconstruction quality. The method’s video generation ability is evaluated using Fréchet Video Distance (FVD). Appendix B holds more details.
5.2 Main Results
Generation Quality. In Table 2, generation quality is evaluated in two ways. First, 3D scene quality is assessed by FID of rendered views (last row of Table 2), with qualitative comparisons in Figure 5. Lacking a quantitative baseline, we compare our pipeline’s renderings with 3DGS. The last two rows of Table 2 show that our enhanced FTGS improves visual quality, especially in novel views. Second, video quality is assessed using extended 16-frame MagicDrive [11] as a baseline. MagicDrive3D significantly boosts video quality (shown by FVD), proving the effectiveness of the proposed relative camera pose embedding for enhancing temporal quality. More qualitative comparisons can be found in Appendix G.
(a) Original view.
(b) Move forward 0.1m.
(c) Move forward 1m.
(d) Move forward 1.6m.
Figure 7: Application of rendering object-level dynamics. After scene generation, we can segment and move the vehicle (the one on the right) in 3D to render a dynamic object.
Controllability. MagicDrive3D accepts 3D bounding boxes, BEV map, and text as control signals, each of which possesses the capacity to independently manipulate the scene. Figure 6 shows such controllability, where we edit a scene from the nuScenes validation set. Clearly, MagicDrive3D effectively alters the scene generation to align with various control signals.
Furthermore, Section 5.3 shows that the rendered views of the generated scenes can be used as training data to improve the performance of perception models. Such results can serve as further evidence of the controllability.
Setting Method no rig depth+0.5m pitch-5 yaw+5 yaw-5
vehicle only real data 17.14 16.63 15.50 16.99 15.94
road only real data 54.94 54.56 53.82 54.20 53.67
5.3 Data Augmentation for Perception Tasks
We demonstrate that street 3D scene generation can serve as a data engine for perception tasks, leveraging the advantage of any-view rendering ability of our method to improve viewpoint robustness for perception models [15]. We employ CVT [46] and the BEV segmentation task following the evaluation protocols of Zhou and Krähenbühl [46], Gao et al. [11]. By incorporating 4 different rigs on the FRONT camera and adding rendered views for training, the negative impact from viewpoint changes is alleviated (Table 3), exemplifying the utility of street scene generation in training perception tasks.
@@ 5.3 Data Augmentation for Perception Tasks
train
view
vary-t 3DGS 0.0189 30.1191 0.9261 0.1259
@@
test
view
vary-t 3DGS 0.0890 17.9879 0.4378 0.4648
@@
Table 4: Performance comparison between the enhanced GS and its baseline. We randomly sample 100 scenes from the nuScenes val. set for evaluation. “cc” refers to color correction from [1]. Although 3DGS does not consider appearance differences, we apply “cc” to it for fair comparisons.
5.4 Application: Render Object-level Dynamic
MagicDrive3D is capable of generating 3DGS representations of scenes, which facilitates various applications in scene editing. Our method utilizes metric scale modeling, guaranteeing that any modifications made to the scenes accurately reflect real-world physical distances. As illustrated in Figure 7, we segmented the generated GS and repositioned the object on the right. The resulting scene GS supports rendering effectively.
@@ 5.4 Application: Render Object-level Dynamic
Table 5: Ablation study on each component of the enhanced GS, with settings in Table I (appendix) and 100 random scenes from the nuScenes validation set.